A positron is a type of antimatter that is just like an electron collides with an electron, their masses are completely annihilated (combined) and converted to energy in the form of gamma rays. (both particles have a rest mass of 9.11 x 10-31 kg. Determine the energy produce.


Hint:


The mass of positron and electron combined during collision, therefore m(positron + electron) Speed of light (3.0x 10 m/s)

With solution


1
Expert's answer
2021-05-21T10:43:11-0400

"E=mc^2=9.11\cdot10^{-31}\cdot(3\cdot10^8)^2=8.2\cdot10^{-14}\ (J)=512\ (keV)"


"E_T=2E=2\cdot512=1025\ (keV)" . Answer